# Reducing quantum error correction overhead using soft information

### v1

@@ I Introduction
Improvements in classical decoders can offer a shortcut to stronger error suppression, but the accuracy of the decoder is often limited by how much information it has about the underlying system [8]. In standard QEC schemes such as surface code quantum computing [9, 4], successive rounds of stabiliser measurements are performed to track the state of a logical qubit against potential errors. This measurement data is typically converted from an analogue to a digital representation and classified to a binary measurement outcome before being passed to a classical decoder [10]. However, the binary classification process is sub-optimal as it throws away valuable information about the measurement process that can be used to improve decoding accuracy [11]. By using a probability estimator that captures the likelihood of a given measurement signal originating from state |$\ket{0}$ or |$\ket{1}$, denoted as soft information, the decoder can more accurately track measurement errors over the course of a QEC protocol, enhancing the code threshold and making the decoder more resilient to classification errors.
@@
We see that in simulated quantum memory experiments, soft decoders provide consistently higher $\Lambda$-factors than hard decoders for both superconducting and neutral atom platforms. In noise regimes where measurement classification errors are unlikely, soft decoding offers next to no advantage, but in more realistic settings where measurement classification is a lower-fidelity operation than gates, soft decoding offers between $10\%$- and $20\%$-improvements in $\Lambda$. Our modelling shows that the scaling advantages of soft decoding are maintained even when measurement times are shorter than ideal—offering faster QEC cycles without a penalty in the error suppression performance compared to hard decoders. For simulated bivariate bicycle codes, we find soft information to reduce logical error rates by an order of magnitude, showcasing that the benefits of the technique are applicable across a wide range of QEC codes.
The paper is structured as follows. A summary of the soft measurement model for superconducting and neutral atom platforms is described in section II.1, and our chosen figures of merit are described in section II.2. In section II.3, we study the impact of soft decoding for surface code experiments on both superconducting and neutral atom platforms, leveraging soft and hard variants of LCD as our decoder. In section II.3.2, we simulate error suppression against the readout duration of superconducting and neutral atom platforms, likewise decoding the results with soft and hard LCD. In section II.4 we show how soft-BP decoders benefit from soft measurement information, benchmarking the decoders on BB codes on a simulated neutral atom platform. Section III concludes the work and provides pointers on effectively incorporating soft information into the decoding workflows of future QEC experiments.
@@ II.1 Soft measurement model
$\begin{split}{p_{\textrm{S}}}_{\mid_{\bar{\mu}=0}}&=\int_{f^{(\bar{\mu}=1)}>f^{(\bar{\mu}=0)}}f^{(\bar{\mu}=1)}(\mu)\text{d}\mu\\ {p_{\textrm{S}}}_{\mid_{\bar{\mu}=1}}&=\int_{f^{(\bar{\mu}=0)}>f^{(\bar{\mu}=1)}}f^{(\bar{\mu}=0)}(\mu)\text{d}\mu\end{split}$ (3)
where the measurement distributions $f^{(\bar{\mu}=0)}(\mu)$ and $f^{(\bar{\mu}=1)}(\mu)$ depend on the measurement time $\tau_{M}$, among other qubit-specific parameters. To capture the frequency of soft errors in our models, we use the average soft flip probability $p_{\textrm{S}}\!=\!({p_{\textrm{S}}}_{\mid_{\bar{\mu}=0}}+{p_{\textrm{S}}}_{\mid_{\bar{\mu}=1}})/2$.
When binary measurements are used, the duration of the measurement operation is typically optimised for maximum measurement fidelity $1-p_{\textrm{S}}$ [29, 30], but simulations show that a shorter readout duration can result in a higher logical fidelity when QEC is applied thanks to a reduction in idling errors during measurement [11]. In our model, used in section II.3 and section II.4, we use measurement times $\tau_{M}$ and error probabilities according to current state-of-the art devices from Google Quantum AI [1] for superconducting devices, and a $Z$-biased noise model along with an appropriate measurement model for neutral atom based devices [31, 23].
@@ II.2 Figures of merit
To evaluate $\Lambda$, we generate instances of rotated planar codes with varying code distances $d=\{5,7,9\}$ and simulate the logical fidelity of a quantum memory experiment taking $T$ syndrome extraction rounds. By decoding the resulting syndrome measurement data, we compare the state of the logical qubit at the end of the experiment with decoder corrections applied against the known state of the logical qubit. By repeating the experiment over multiple shots, we obtain the logical error probability $P_{L}(d,T)$ for each code distance $d$ and number of rounds $T$. To extract $\Lambda$ from this data, we first use the relation
$P_{L}(d,T)=\frac{1}{2}\left[1-(1-2\epsilon_{L}(d))^{T}\right]$ (5)
to compute the logical error rate per round $\epsilon_{L}(d)$, and then use a linear fit of $\log(\epsilon_{L})$ to compute $\Lambda$ the details of this process are given in section IV.2.
In the case of bivariate bicycle codes, we do not extract a $\Lambda$-value, as we would need codes that share the same defining bivariate polynomials. The BB codes reported in Ref. [19] were found by a numerical search and do not all belong to the same family. Hence, to study the impact of soft information on these codes we consider the logical error probability $P_{L}$ as the figure of merit.
@@ II.3.2 Measurement time optimisation
Mid-circuit measurements are integral to the practical realisation of many QEC codes such as the surface code, but the act of measurement is often among the most error-prone operations in current QEC experiments [10, 1]. Although it is often possible to improve measurement fidelity by setting longer measurement times, this comes at the cost of increased idling errors and a slower logical clock speed due to the additional time needed for syndrome extraction. To mitigate these challenges, soft information decoding promises improved resilience against measurement errors—making it possible to shorten the measurement time in an experiment without incurring a penalty in the logical fidelity. In this section, we study the impact of the measurement time $\tau_{M}$ on the error suppression rate $\Lambda$ by simulating quantum memory experiments on superconducting and neutral atom qubit platforms.
For superconducting qubits, we take the SI1000 noise model as our baseline circuit-level noise model and introduce a variation to its error channel probabilities that depends on measurement and gate durations, as described in section IV.1.1. In our parametrisation, the measurement time $\tau_{M}$ alters the probability of depolarising errors while qubits are idling, being measured or reset. The device operation times are set as $\tau_{1q}=20~{}\text{ns}$ and $\tau_{2q}=40~{}\text{ns}$ for one-qubit and two-qubit gates respectively and $\tau_{R}=20~{}\text{ns}$ for resets. The soft measurement response function is likewise dependent on $\tau_{M}$, with a signal-to-noise ratio (SNR) that is proportional to $\tau_{M}$ for measurement times $\tau_{M}\ll\textrm{T}_{1}$. Due to amplitude damping with $\textrm{T}_{1}=100~{}\mu\textrm{s}$, in line with the current state-of-the art devices [1], qubits experience time-dependent decay both while idling and during measurement [36]. For neutral atom qubits, the dominant source of error is the two-qubit error rate—with idle noise being insignificant. We use the same $Z$-biased circuit-level noise model (see section IV.1.2) as done previously, but the measurement response function is varied according to $\tau_{M}$.
@@
Figure 4: LCD: error suppression rate $\Lambda$ versus measurement time $\tau_{M}$. $\Lambda$ is obtained from simulations of rotated planar code quantum memory experiments with $T=10$ rounds of syndrome extraction. In panel a) we plot $\Lambda$ for superconducting qubits under physical error rate $p=0.3\%$, classification error rate $p_{\textrm{S}}=5p$ at $\tau_{M}=500~{}\textrm{ns}$ and $\textrm{T}_{1}=100~{}\mu\textrm{s}$. In panel b), we plot $\Lambda$ for neutral atom qubits, with $p=1\%$ and $Z$-bias $100$. The soft flip probability $p_{\textrm{S}}$ varies based on $\tau_{M}$. On both qubit platforms, a higher $\Lambda$ is reached with soft compared to hard decoding. Error bars in each plot correspond to hypotheses with a likelihood within a factor of 10 of the maximum likelihood hypothesis, given the sampled data.
As the measurement time $\tau_{M}$ dictates the duration of a QEC cycle, shortening measurements without losing accuracy is a key challenge in the design of error-corrected quantum computers. Taking the maximum $\Lambda$ for the hard decoder as a benchmark, Fig. 4 shows that measurement duration can be shortened by $55\%$ for superconducting qubits and by $40\%$ for neutral atoms without incurring a loss in error suppression performance. At shorter-than-ideal measurement times, the soft decoder shows much higher $\Lambda$-values than the hard decoder, owing to its better resilience against measurement classification errors. As the superconducting noise model has measurement-time-dependent idling errors, we find that the optimal $\tau_{M}$ for achieving maximum $\Lambda$ across $X$-and $Z$-basis is $35\%$ shorter for the soft decoder than the hard decoder. The difference in $\Lambda$ between the soft and the hard decoder vanishes for long measurement durations, as the improved SNR reduces the frequency of classification errors in the system.
In the neutral atom picture, time-dependent errors during idling are not significant thanks to the long qubit coherence times and fast single-qubit and two-qubit gates [37, 31]. When varying the measurement time, only the classification error probability is affected, leading to both the soft and the hard decoders achieving maximum $\Lambda$ at $\tau_{M}\approx 150~{}\mu\textrm{s}$ as seen in Fig. 4 b). Notably, the improvement in $\Lambda$ for the soft decoder compared to the hard decoder is sustained even for $\tau_{M}>200~{}\mu\textrm{s}$. This is explained by the presence of background noise and decay-related classification errors in the system that persist despite the long measurement time.
@@ II.4.1 Soft information in BP-Matching
Figure 5: BM on superconducting qubits: error suppression rate $\Lambda$ versus physical qubit fidelity $\mathcal{F}=1-p$. Data from a simulated quantum memory experiment with $T=10$ rounds of syndrome extraction and $N=5\times 10^{4}$ shots per data point. In panel a), we show $\Lambda$ for a noise regime where soft flips are sub-dominant $p_{\textrm{S}}/p=1$, and in panel b) we show $\Lambda$ when soft flips are a significant component of the error model $p_{\textrm{S}}/p=5$. Error bars in each plot correspond to hypotheses with a likelihood within a factor of 10 of the maximum likelihood hypothesis, given the sampled data.
@@ II.4.2 Soft information in BP+OSD
Bivariate bicycle (BB) codes are quantum low-density parity check (LDPC) codes that exploit long-range qubit connectivity to achieve higher encoding rates than comparably sized surface codes, while maintaining a high error threshold [19]. This high connectivity needed for their implementation can be easily realised via qubit shuttling on atomic qubits [41], making neutral atoms a viable candidate for the realisation of these codes in the near term. The Gross code $[[n,k,d]]\!\!=\!\![[144,12,12]]]$ is among the most promising for near-term demonstrations because of its high encoding rate. In this section, we analyse the performance of BB codes under a circuit-level noise model for neutral atom qubits, decoding the resulting soft measurements with a BP decoder augmented with soft information. To break the degeneracy problem otherwise inherent in the BP algorithm, we combine soft-BP with Ordered Statistics Decoding (OSD) [42, 43, 40] as a post-processing step. We note that significantly faster decoders have been proposed [44, 45], however these were not available to us when these simulations were conducted.
@@
To evaluate error suppression performance, we take the syndrome extraction circuits in [46] for two instances of BB codes $[[n,k,d]]\!\!=\!\![[72,12,6]]$ and $[[144,12,12]]$, and replace the noise with our circuit-level neutral atom noise model, shown in section IV.1.2. As before, we evaluate decoder performance under two noise regimes—LF with $p_{\textrm{S}}=p$ and HF with $p_{\textrm{S}}=5p$, decoding with both a soft-information-aware BP+OSD decoder and a traditional hard BP+OSD decoder. We set up BP with $100$ BP iterations, min-sum as the BP method of choice, combination sweep as the OSD method and an OSD order of 00. The measurement error probabilities obtained from soft information are used to dynamically update the error channel probabilities of the BP+OSD decoder using the update channel probabilities function in [47]. As our figure of merit we use the logical error rate $P_{L}$, as we only used two codes with $d=6$ and $d=12$ for this simulation.
@@
In Fig. 6, we evaluate decoding accuracy for BP+OSD under the LF and HF soft-measurement flip regimes. We find that in the LF-regime, soft information provides more than a $90\%$ reduction in the logical error probability across the physical error rate regime examined compared to hard BP+OSD. In the HF-regime, soft information provides more than a $95\%$ reduction in logical errors compared to hard BP+OSD. Our results clearly show that soft information in BP+OSD significantly enhances the error suppression ability of BB codes, especially at the high physical error rates likely to be seen in near-term devices.
@@
Figure 6: Logical error probabilities as a function of physical error rates (p) for bivariate bicycle (BB) codes. We plot logical error probabilities for BB codes with circuit level noise for neutral atom qubits using the BP+OSD decoder. We consider two instances of BB codes: $d\!\!=\!\!6$ for $([[72,12,6]])$ code and the Gross code, $[[144,12,12]]$, with $d\!\!=\!\!12$ We compare the impact of soft information when a) soft measurement flip probability $p_{\textrm{S}}\!\!=\!\!p$. (b) soft measurement flip probability $p_{\textrm{S}}\!\!=\!\!5p$. Error bars in each plot correspond to hypotheses with a likelihood within a factor of 10 of the maximum likelihood hypothesis, given the sampled data.

### v2

@@ I Introduction
Improvements in classical decoders can offer a shortcut to stronger error suppression, but the accuracy of the decoder is often limited by how much information it has about the underlying system [8]. In standard QEC schemes such as surface code quantum computing [9, 4], successive rounds of stabiliser measurements are performed to track the state of a logical qubit against potential errors. This measurement data is typically converted from an analogue to a digital representation and classified to a binary measurement outcome before being passed to a classical decoder [10]. However, the binary classification process is sub-optimal as it throws away valuable information about the measurement process that can be used to improve decoding accuracy [11, 12]. By using a probability estimator that captures the likelihood of a given measurement signal originating from state |$\ket{0}$ or |$\ket{1}$, denoted as soft information, the decoder can more accurately track measurement errors over the course of a QEC protocol, enhancing the code threshold and making the decoder more resilient to classification errors.
@@
We see that in simulated quantum memory experiments, soft decoders provide consistently higher $\Lambda$-factors than hard decoders for both superconducting and neutral atom platforms. In noise regimes where measurement classification errors are unlikely compared to data errors, soft decoding offers next to no advantage, but in more realistic settings where measurement classification is a lower-fidelity operation than gates, soft decoding offers between $10\%$- and $20\%$-improvements in $\Lambda$. Our modelling shows that the scaling advantages of soft decoding are maintained even when measurement times are shorter than ideal—offering faster QEC cycles without a penalty in the error suppression performance compared to hard decoders. For simulated bivariate bicycle codes, we find soft information to reduce logical error probabilities by an order of magnitude, showcasing that the benefits of the technique are applicable across a wide range of QEC codes.
The paper is structured as follows. A summary of the soft measurement model for superconducting and neutral atom platforms is described in section II.1, and our chosen figures of merit are described in section II.2. In section II.3, we study the impact of soft decoding for surface code experiments on both superconducting and neutral atom platforms, leveraging soft and hard variants of LCD as our decoder. In section II.3.2, we simulate error suppression against the readout duration of superconducting and neutral atom platforms, likewise decoding the results with soft and hard LCD. In section II.4 we show how soft-BP decoders benefit from soft measurement information, benchmarking the decoders on BB codes on simulated neutral atom and superconducting platforms. Section III concludes the work and provides pointers on effectively incorporating soft information into the decoding workflows of future QEC experiments.
@@ II.1 Soft measurement model
$\begin{split}{p_{\textrm{S}}}_{\mid_{\bar{\mu}=0}}&=\int_{f^{(\bar{\mu}=1)}>f^{(\bar{\mu}=0)}}f^{(\bar{\mu}=0)}(\mu)\text{d}\mu\\ {p_{\textrm{S}}}_{\mid_{\bar{\mu}=1}}&=\int_{f^{(\bar{\mu}=0)}>f^{(\bar{\mu}=1)}}f^{(\bar{\mu}=1)}(\mu)\text{d}\mu\end{split}$ (3)
where the measurement distributions $f^{(\bar{\mu}=0)}(\mu)$ and $f^{(\bar{\mu}=1)}(\mu)$ depend on the measurement time $\tau_{\textrm{M}}$, among other qubit-specific parameters. To capture the frequency of soft errors in our models, we use the average soft flip probability $p_{\textrm{S}}\!=\!({p_{\textrm{S}}}_{\mid_{\bar{\mu}=0}}+{p_{\textrm{S}}}_{\mid_{\bar{\mu}=1}})/2$.
When binary measurements are used, the duration of the measurement operation is typically optimised for maximum measurement fidelity $1-p_{\textrm{S}}$ [30, 31], but simulations show that a shorter readout duration can result in a higher logical fidelity when QEC is applied thanks to a reduction in idling errors during measurement [12]. In our model, used in section II.3 and section II.4, we use measurement times $\tau_{\textrm{M}}$ and error probabilities according to current state-of-the art devices from Google Quantum AI [1] for superconducting devices, and a $Z$-biased noise model along with an appropriate measurement model for neutral atom based devices [32, 24].
@@ II.2 Figures of merit
To evaluate $\Lambda$, we generate instances of rotated planar codes with varying code distances $d=\{5,7,9\}$ and simulate the logical fidelity of a quantum memory experiment taking $T$ syndrome extraction rounds. By decoding the resulting syndrome measurement data, we compare the state of the logical qubit at the end of the experiment with decoder corrections applied against the known state of the logical qubit. By repeating the experiment over multiple shots, we obtain the logical error probability $P_{\textrm{L}}(d,T)$ for each code distance $d$ and number of rounds $T$. To extract $\Lambda$ from this data, we first use the relation
$P_{\textrm{L}}(d,T)=\frac{1}{2}\left[1-(1-2\epsilon_{\textrm{L}}(d))^{T}\right]$ (5)
to compute the logical error rate per round $\epsilon_{\textrm{L}}(d)$, and then use a linear fit of $\log(\epsilon_{\textrm{L}})$ to compute $\Lambda$ the details of this process are given in section IV.2.
In the case of bivariate bicycle codes, we do not extract a $\Lambda$-value, as we would need codes that share the same defining bivariate polynomials. The BB codes reported in Ref. [20] were found by a numerical search and do not all belong to the same family. Hence, to study the impact of soft information on these codes we consider the logical error probability $P_{\textrm{L}}$ as the figure of merit.
@@ II.3.2 Measurement time optimisation
Mid-circuit measurements are integral to the practical realisation of many QEC codes such as the surface code, but the act of measurement is often among the most error-prone operations in current QEC experiments [10, 1]. Although it is often possible to improve measurement fidelity by setting longer measurement times, this comes at the cost of increased idling errors and a slower logical clock speed due to the additional time needed for syndrome extraction. To mitigate these challenges, soft information decoding promises improved resilience against measurement errors—making it possible to shorten the measurement time in an experiment without incurring a penalty in the logical fidelity. In this section, we study the impact of the measurement time $\tau_{\textrm{M}}$ on the error suppression rate $\Lambda$ by simulating quantum memory experiments on superconducting and neutral atom qubit platforms.
For superconducting qubits, we take the SI1000 noise model as our baseline circuit-level noise model and introduce a variation to its error channel probabilities that depends on measurement and gate durations, as described in section IV.1.1. In our parametrisation, the measurement time $\tau_{\textrm{M}}$ alters the probability of depolarising errors while qubits are idling, being measured or reset. The device operation times are set as $\tau_{\textrm{1q}}=20~{}\text{ns}$ and $\tau_{\textrm{2q}}=40~{}\text{ns}$ for one-qubit and two-qubit gates respectively and $\tau_{\textrm{R}}=20~{}\text{ns}$ for resets. The soft measurement response function is likewise dependent on $\tau_{\textrm{M}}$, with a signal-to-noise ratio (SNR) that is proportional to $\tau_{\textrm{M}}$ for measurement times $\tau_{\textrm{M}}\ll\textrm{T}_{1}$. Due to amplitude damping with $\textrm{T}_{1}=100~{}\mu\textrm{s}$, in line with the current state-of-the art devices [1], qubits experience time-dependent decay both while idling and during measurement [37]. For neutral atom qubits, the dominant source of error is the two-qubit error rate—with idle noise being insignificant. We use the same $Z$-biased circuit-level noise model (see section IV.1.2) as done previously, but the measurement response function is varied according to $\tau_{\textrm{M}}$.
@@
Figure 4: LCD: error suppression rate $\Lambda$ versus measurement time $\tau_{\textrm{M}}$. $\Lambda$ is obtained from simulations of rotated planar code quantum memory experiments with $T=10$ rounds of syndrome extraction. In panel a) we plot $\Lambda$ for superconducting qubits under physical error rate $p=0.3\%$, classification error rate $p_{\textrm{S}}=5p$ at $\tau_{\textrm{M}}=500~{}\textrm{ns}$ and $\textrm{T}_{1}=100~{}\mu\textrm{s}$. In panel b), we plot $\Lambda$ for neutral atom qubits, with $p=1\%$ and $Z$-bias $100$. The soft flip probability $p_{\textrm{S}}$ varies based on $\tau_{\textrm{M}}$. On both qubit platforms, a higher $\Lambda$ is reached with soft compared to hard decoding. Error bars in each plot correspond to hypotheses with a likelihood within a factor of 10 of the maximum likelihood hypothesis, given the sampled data.
As the measurement time $\tau_{\textrm{M}}$ dictates the duration of a QEC cycle, shortening measurements without losing accuracy is a key challenge in the design of error-corrected quantum computers. Taking the maximum $\Lambda$ for the hard decoder as a benchmark, Fig. 4 shows that measurement duration can be shortened by $55\%$ for superconducting qubits and by $40\%$ for neutral atoms without incurring a loss in error suppression performance. At shorter-than-ideal measurement times, the soft decoder shows much higher $\Lambda$-values than the hard decoder, owing to its better resilience against measurement classification errors. As the superconducting noise model has measurement-time-dependent idling errors, we find that the optimal $\tau_{\textrm{M}}$ for achieving maximum $\Lambda$ across $X$-and $Z$-basis is $35\%$ shorter for the soft decoder than the hard decoder. The difference in $\Lambda$ between the soft and the hard decoder vanishes for long measurement durations, as the improved SNR reduces the frequency of classification errors in the system.
In the neutral atom picture, time-dependent errors during idling are not significant thanks to the long qubit coherence times and fast single-qubit and two-qubit gates [38, 32]. When varying the measurement time, only the classification error probability is affected, leading to both the soft and the hard decoders achieving maximum $\Lambda$ at $\tau_{\textrm{M}}\approx 150~{}\mu\textrm{s}$ as seen in Fig. 4 b). Notably, the improvement in $\Lambda$ for the soft decoder compared to the hard decoder is sustained even for $\tau_{\textrm{M}}>200~{}\mu\textrm{s}$. This is explained by the presence of background noise and decay-related classification errors in the system that persist despite the long measurement time.
@@ II.4.1 Soft information in BP-Matching
Figure 5: BM on superconducting qubits: error suppression rate $\Lambda$ versus physical qubit fidelity $\mathcal{F}=1-p$. Data from a simulated quantum memory experiment with $T=10$ rounds of syndrome extraction and $N=5\times 10^{5}$ shots per data point. In panel a), we show $\Lambda$ for a noise regime where soft flips are sub-dominant $p_{\textrm{S}}/p=1$, and in panel b) we show $\Lambda$ when soft flips are a significant component of the error model $p_{\textrm{S}}/p=5$. Error bars in each plot correspond to hypotheses with a likelihood within a factor of 10 of the maximum likelihood hypothesis, given the sampled data.
@@ II.4.2 Soft information in BP+OSD
Bivariate bicycle (BB) codes are quantum low-density parity check (LDPC) codes that exploit long-range qubit connectivity to achieve higher encoding rates than comparably sized surface codes, while maintaining a high error threshold [20]. This high connectivity needed for their implementation can be easily realised via qubit shuttling on atomic qubits [42], making neutral atoms a viable candidate for the realisation of these codes in the near term. Alternatively, the degree-6 Tanner graph of the BB codes can be decomposed into two planar subgraphs facilitating their practical implementation on a superconducting qubit architecture.
The Gross code $[[n,k,d]]\!\!=\!\![[144,12,12]]]$ is among the most promising for near-term demonstrations thanks to its high encoding rate. In this section, we analyse the performance of BB codes under a circuit-level noise model for neutral atom qubits as well as the superconducting qubits, decoding the resulting soft measurements with a BP decoder augmented with soft information. To break the degeneracy problem otherwise inherent in the BP algorithm, we combine soft-BP with Ordered Statistics Decoding (OSD) [43, 44, 41] as a post-processing step. We note that significantly faster decoders have been proposed [45, 46], however these were not available to us when these simulations were conducted.
@@
To evaluate error suppression performance, we take the syndrome extraction circuits in [47] for two instances of BB codes $[[n,k,d]]\!\!=\!\![[72,12,6]]$ and $[[144,12,12]]$, and replace the noise in these circuits with an architecture-specific circuit-level noise model. For the neutral atom simulations, we use a $Z$-biased noise model as described in section IV.1.2, while the superconducting simulations have a SI1000 noise model applied, as detailed in section IV.1.1. As before, we evaluate decoder performance under two noise regimes—LF with $p_{\textrm{S}}=p$ and HF with $p_{\textrm{S}}=5p$, decoding with both a soft-information-aware BP+OSD decoder and a traditional hard BP+OSD decoder. We set up BP with $100$ BP iterations, min-sum as the BP method of choice, combination sweep as the OSD method and an OSD order of 00. The measurement error probabilities obtained from soft information are used to dynamically update the error channel probabilities of the BP+OSD decoder using the update channel probabilities function in [48]. As our figure of merit we use the logical error probability ($P_{\textrm{L}}$) instead of $\Lambda$, as we only used two codes with $d=6$ and $d=12$ for this simulation.
@@
In Fig. 6, we evaluate decoding accuracy for BP+OSD under the LF and HF soft-measurement flip regimes for both superconducting and neutral atom qubit platforms. We find that in the LF-regime, soft information provides more than a $90\%$ reduction in the logical error probability across the physical error rate regime examined compared to hard BP+OSD. In the HF-regime, soft information provides more than a $95\%$ reduction in logical errors compared to hard BP+OSD. Our results clearly show that soft information in BP+OSD significantly enhances the error suppression ability of BB codes, especially at the high physical error rates likely to be seen in near-term devices.
@@
Figure 6: Logical error probability $P_{\textrm{L}}$ as a function of the physical error rate $p$ for bivariate bicycle (BB) codes decoded with BP+OSD. We consider two instances of BB codes: $d\!\!=\!\!6$ for the $[[72,12,6]]$ code and $d\!\!=\!\!12$ for the $[[144,12,12]]$ Gross code. In panels a) and b) we plot the logical error probability $P_{\textrm{L}}$ for simulated neutral atom qubits under two different regimes of soft measurement error probability: $p_{\textrm{S}}/p=1$ for sub-dominant soft measurement errors, and $p_{\textrm{S}}/p=5$ for dominant soft measurement errors. Each data point in panels a) and b) uses $5\times 10^{4}$ shots. In panels c) and d) we plot the logical error probability $P_{\textrm{L}}$ for simulated superconducting qubits using $5\times 10^{5}$ shots per data point, with $p_{\textrm{S}}/p=1$ and $p_{\textrm{S}}/p=5$ in panels c) and d) respectively. Error bars in all the plots correspond to hypotheses with a likelihood within a factor of 10 of the maximum likelihood hypothesis, given the sampled data.